Given the system of linear equations,(3x-4y=5 2x+3y=8)
Re-write both equations in the y = mx + b form.
What is the slope of the first equation? ………………..
What is the slope of the second equation? ……………
Based on the slopes, will these lines intersect or they are parallel?
……………………………………………………………………………..

Answer by lwsshak3(11628)

Given the system of linear equations,(3x-4y=5 2x+3y=8)
Re-write both equations in the y = mx + b form.
What is the slope of the first equation? ………………..
What is the slope of the second equation? ……………
Based on the slopes, will these lines intersect or they are parallel?
**
First equation:
3x-4y=5
4y=3x-5
y=(3/4)x-5/4
slope=3/4
..
Second equation:
2x+3y=8
3y=-2x+8
y=(-2/3)x+8/3
slope=-2/3
..
lines will intersect